### v1

@@ 3 Discrete Elastic Rods Preliminaries
Here, we define $E_{\mathrm{stretch}}(\mathbf{x})$ with $\mathbf{x}$ alone because stretching is independent of $\bm{\theta}$. While velocity-based damping can also be incorporated into the optimization (1) [George2018], it is irrelevant to our rest shape optimization since we focus on the static equilibrium case. We can obtain the generalized positions for the next time step by solving the optimization problem (1) with a gradient-based optimizer (e.g., Newton’s method [NoceWrig06]) using the gradient of the DER objective with respect to $\mathbf{q}$, $\nabla_{\mathbf{q}}E_{\mathrm{DER}}(\mathbf{q})$, while handling the fixed variables as Dirichlet boundary conditions. We note that the gradient with respect to the fixed variables is not defined (or, in practice, it can be assumed 0 for convenience). In our framework, we use a single Newton iteration per simulation step for efficiency [Kim2022deformables, Huang2023hair] (although this approach sacrifices stability with larger parameter values [Shi2023curls]). In the following sections, we define each objective term and its gradient.
@@ 3.1 Inertia
$E_{\mathrm{inertia}}(\mathbf{q})=\frac{1}{2\Delta t^{2}}\left\lVert\mathbf{q}-\mathbf{q}^{*}\right\rVert^{2}_{\mathbf{M}},\quad\mathbf{q}^{*}=\mathbf{q}^{t}+\Delta t\mathbf{\dot{q}}^{t}+\Delta t^{2}\mathbf{M}^{-1}\mathbf{f}_{\mathrm{ext}},$ (3)
where $\mathbf{q}^{*}$ denotes the predicted generalized positions, $\mathbf{q}^{t}$ and $\mathbf{\dot{q}}^{t}$ denote the generalized positions and velocities at time $t$, respectively, $\Delta t\ (>0)$ time step size, $\mathbf{M}\in\mathbb{R}^{(4N-1)\times(4N-1)}$ a diagonal, generalized mass matrix [Bergou:2010:DVT:1778765.1778853, jawed2018primer], and $\mathbf{f}_{\mathrm{ext}}\in\mathbb{R}^{4N-1}$ a generalized external force vector. While sag-free simulation typically assumes gravity as the sole external force, we can nevertheless easily support other constant external forces, such as wind forces or load on specific vertices [Perez2015rod] or edges.
@@ 3.2 Stretching
where $c_{\mathrm{stretch}}$ denotes the stiffness coefficient for stretching, and $\mathbf{l}_{i}\ (=\left\lVert\mathbf{x}_{i+1}-\mathbf{x}_{i}\right\rVert_{2})$ denotes the length of edge $i$. (Unlike the work of [Bergou:2010:DVT:1778765.1778853], which uses Young’s modulus and shear modulus, we use different coefficients for each energy to individually control their strengths.) Given that the vertices $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ are fixed, there is no stretching energy contribution $E_{\mathrm{stretch},0}(\mathbf{x}_{0},\mathbf{x}_{1})$ from edge 00 (the first edge at the root end).
@@ 3.3 Bending
We define the bending energy on vertex $i$ with the four-dimensional curvature $\bm{\kappa}_{i}\in\mathbb{R}^{4}$ (i.e., $\bm{\kappa}=[\bm{\kappa}_{1}^{T},\ldots,\bm{\kappa}_{N-2}^{T}]^{T}\in\mathbb{R}^{4N-8}$) and its corresponding rest curvature $\bm{\bar{\kappa}}_{i}\in\mathbb{R}^{4}$ in 3D [Bergou2008, Fei2019] as
@@
where $c_{\mathrm{bend}}$ denotes the stiffness coefficient for bending energy. Here, we do not use the bending model with the two-dimensional curvature (and rest curvature) via the averaged material frames from [Bergou:2010:DVT:1778765.1778853] because their model would fail to correctly evaluate the bending of strands [Gornowicz2015, Panetta2019]. In addition, the bending model (12) [Bergou2008, Fei2019] is slightly simpler than the one presented by [Panetta2019] (which modifies the rest length term) and avoids introducing additional complexity when differentiating (12) with respect to the rest length (see Sec. 4.1.3).
@@ 3.4 Twisting
where the twisting energy defined for vertex $i$ is given with twist $\mathbf{m}_{i}$ ($\mathbf{m}=[\mathbf{m}_{1},\ldots,\mathbf{m}_{N-2}]^{T}\in\mathbb{R}^{N-2}$), which takes the reference twist into account [Bergou:2010:DVT:1778765.1778853, Kaldor2010, jawed2018primer, Panetta2019], and its corresponding rest twist $\mathbf{\bar{m}}_{i}$ by
$E_{\mathrm{twist},i}(\mathbf{x}_{i-1},\bm{\theta}_{i-1},\mathbf{x}_{i},\bm{\theta}_{i},\mathbf{x}_{i+1})=\frac{1}{2}\left(\frac{c_{\mathrm{twist}}\pi r^{4}}{\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i}}\right)(\mathbf{m}_{i}-\mathbf{\bar{m}}_{i})^{2},$ (16)
@@
$\nabla E_{\mathrm{twist},i}(\mathbf{x}_{i-1},\bm{\theta}_{i-1},\mathbf{x}_{i},\bm{\theta}_{i},\mathbf{x}_{i+1})=\left(\frac{c_{\mathrm{twist}}\pi r^{4}}{\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i}}\right)(\mathbf{m}_{i}-\mathbf{\bar{m}}_{i})\nabla\mathbf{m}_{i}.$ (17)
@@ 4 Rest Shape Optimization
We define the rest shape parameter $\mathbf{\bar{s}}=[\mathbf{\bar{l}}_{1},\ldots,\mathbf{\bar{l}}_{N-2},\bm{\bar{\kappa}}^{T},\mathbf{\bar{m}}^{T}]^{T}\in\mathbb{R}^{6N-12}$ since $(6N-12)=(N-2)+4(N-2)+(N-2)$. Note that $\mathbf{\bar{s}}$ excludes $\mathbf{\bar{l}}_{0}$ because the first two root-end vertices are fixed, and thus the edge between them should also be fixed. We aim to optimize $\mathbf{\bar{s}}$ based on minimization of the kinetic energy with a regularizer subject to box constraints. Specifically, we formulate the rest shape optimization as
@@
where $\mathbf{\bar{s}_{\mathrm{min}}}\in\mathbb{R}^{6N-12}$ and $\mathbf{\bar{s}_{\mathrm{max}}}\in\mathbb{R}^{6N-12}$ denote the lower and upper bounds imposed on $\mathbf{\bar{s}}$, respectively, and $F_{\mathrm{kin}}(\mathbf{\bar{s}})$ and $F_{\mathrm{reg}}(\mathbf{\bar{s}})$ denote the kinetic energy objective and a regularizer, respectively. Here, we divide $F_{\mathrm{kin}}(\mathbf{\bar{s}})$ and $F_{\mathrm{reg}}(\mathbf{\bar{s}})$ by $\Delta t^{2}$, anticipating that it will cancel with the $\Delta t^{2}$ in (20) and (29), and noting that this scaling by a constant value has no effect on the rest shape optimization (18). In addition, since $\mathbf{\bar{l}}_{0}$ is not in $\mathbf{\bar{s}}$, we do not define the Jacobian with respect to $\mathbf{\bar{l}}_{0}$ (similar to the fixed variables in Sec. 3), and thus $\mathbf{\bar{l}}_{0}$ is not modified by the rest shape optimization (18).
@@ 4.1 Generalized Kinetic Energy Objective
Considering the Jacobian $\mathbf{J}_{\mathrm{DER}}(=\frac{\partial\mathbf{f}_{\mathrm{DER}}}{\partial\mathbf{\bar{s}}})\in\mathbb{R}^{(4N-8)\times(6N-12)}$, the gradient of (20) can be computed as
@@
and the Hessian can be approximated in the Gauss-Newton style by
@@ 4.1.1 Inertia
This result indicates that the inertia force is independent of $\mathbf{\bar{l}}$, and thus the Jacobian $\frac{\partial\mathbf{f}_{\mathrm{inertia}}}{\partial\mathbf{\bar{s}}}=0$. In addition, (23) is also independent of $\Delta t$, and thus the actual value of $\Delta t$ has no effect on the rest shape optimization (18).
@@ 4.1.4 Twisting
4.1.5 Discussion: Kinetic-Energy-Norm vs. Force-Norm
Eliding unnecessary coefficients for a simpler analysis, the objective based on the $L^{2}$ norm of forces is given by $\left\lVert\mathbf{f}\right\rVert_{2}^{2}$ [Twigg:2011:OSS:2019406.2019437] ($\mathbf{f}$: forces) whereas our kinetic-energy-based objective is $\left\lVert\mathbf{f}\right\rVert_{\mathbf{M}^{-1}}^{2}$ due to (20). As an example, consider the inertia force (23) due to gravity $\mathbf{g}$ because the inertia is the unique trigger for non-zero elastic forces in our framework. In this case, $\mathbf{f}=\mathbf{M}\mathbf{g}$, and thus we get $\mathbf{g}^{T}\mathbf{M}^{2}\mathbf{g}$ for the force-norm-based objective versus $\mathbf{g}^{T}\mathbf{M}\mathbf{g}$ for ours. As such, when mass/inertia of the strands is too small or large (typically the inertia is rather small for thin strands [Bergou2008, Bergou:2010:DVT:1778765.1778853], e.g., $10^{-8}\ \mathrm{kg\cdot m^{2}}$), squaring $\mathbf{M}$ can negatively impact the numerical conditioning of the system, requiring extended precision for the optimization to consistently succeed [Twigg:2011:OSS:2019406.2019437]. Our approach neatly avoids this problematic squaring.
@@ 4.2 Regularizer for Rest Shape Parameters
where $\alpha$ denotes a tunable regularizer coefficient (we typically set $\alpha=10^{-5}$). Although the given system may already be overdetermined (having a unique solution) depending on the initial strand geometry, even in such cases, using the regularizer generally has a positive effect in mitigating overfitting. The previous work [Twigg:2011:OSS:2019406.2019437] proposed using a regularizer based on the force norm, but we found it to be unnecessary because the regularizer in our framework is used to ensure that the system is overdetermined, while the box constraints can restrict the changes in the rest shape parameters in a more controlled way (see Sec. 4.3).
@@ 4.3 Box Constraints for Rest Shape Parameters
Given the turning angle $\phi_{i}$ at vertex $i$ [Bergou2008], we can derive the relation between the norm of the curvature and turning angle as $\frac{1}{2}\left\lVert\bm{\kappa}_{i}\right\rVert_{2}^{2}=\left(2\tan\frac{\phi_{i}}{2}\right)^{2}$. Thus, if we have the change of the turning angle $\Delta\phi$ at $\phi_{i}=0$ with the corresponding change of the curvature $\Delta\kappa$ (assuming $\Delta\kappa=\Delta\bm{\kappa}_{i,0}=\Delta\bm{\kappa}_{i,1}=\Delta\bm{\kappa}_{i,2}=\Delta\bm{\kappa}_{i,3}$ for simplicity), we get $\frac{1}{2}(4\Delta\kappa)^{2}=\left(2\tan\frac{\Delta\phi}{2}\right)^{2}$, and thus $|\Delta\kappa|=\sqrt{2}|\tan\frac{\Delta\phi}{2}|$. This relation indicates that, e.g., allowing a maximal change of the turning angle of $\frac{\pi}{2}$ (i.e., $\Delta\phi=\frac{\pi}{2}$) leads to a change of the curvature of $\Delta\kappa=\sqrt{2}$. Based on this analysis for bending, we define $\mathbf{b}$ as
@@ 4.4 Gauss-Newton Solver with Penalty Method
Given the least-squares-style nonlinear objectives arising in our problem, we aim to solve the optimization via the GN method [NoceWrig06]. While enforcing box constraints within Newton-type optimizers is possible [Takahashi2021], it requires accounting for the box constraints when solving inner problems, which leads to an additional computational cost compared to unconstrained linear solves. Considering that our problems are relatively stiff (albeit yielding sparse systems), instead of handling box constraints in the inner problems, we prefer to convert the box-constrained minimization into an unconstrained one via the penalty method [NoceWrig06], thus ensuring unconstrained inner linear systems for greater efficiency.
Specifically, we reformulate the box-constrained minimization (18) into the following unconstrained minimization by converting the box constraints into a penalty objective $F_{\mathrm{box}}(\mathbf{\bar{s}})$ [NoceWrig06]:
$\mathbf{\bar{s}}=\operatorname*{arg\,min}_{\mathbf{\bar{s}}}F(\mathbf{\bar{s}}),\quad F(\mathbf{\bar{s}})=\frac{1}{\Delta t^{2}}\left(F_{\mathrm{kin}}(\mathbf{\bar{s}})+F_{\mathrm{reg}}(\mathbf{\bar{s}})+F_{\mathrm{box}}(\mathbf{\bar{s}})\right),$ (37)
@@
with $\beta$ denoting a penalty parameter (we set $\beta=10^{6}$).
@@
Algorithm 1 shows our rest shape optimization algorithm with the GN solver. To accelerate its convergence via warm starting and to reduce the possibility of falling into suboptimal local minima, we initialize $\mathbf{\bar{s}}$ from the initial $\mathbf{l}$, $\bm{\kappa}$, and $\mathbf{m}$. The inner linear system is guaranteed to be symmetric positive definite (SPD) due to the GN-style Hessian approximation and the regularizer. Given that the system is relatively stiff yet sparse, we solve it using a sparse Cholesky-based direct solver. To ensure a decrease of the objective, we employ a back-tracking line search [NoceWrig06]. In practice, as we enforce the box constraints via the penalty, the rest shape parameters $\mathbf{\bar{s}}$ can slightly violate their bounds. Given these relatively small violations, if necessary, we can project $\mathbf{\bar{s}}$ back into the valid range in each iteration (although this treatment was not needed in any of our examples). We terminate the GN iterations when $\left\lVert\Delta\mathbf{\bar{s}}\right\rVert_{2}$ becomes smaller than a threshold $\epsilon\ (=10^{-5})$. In addition, to ensure that the GN solver ultimately terminates, we halt the iteration if it exceeds $500$ iterations or the back-tracking line search fails.

### v2

@@ 3 Discrete Elastic Rods Preliminaries
Here, we define $E_{\mathrm{stretch}}(\mathbf{x})$ with $\mathbf{x}$ alone because stretching is independent of $\bm{\theta}$. While velocity-based damping can also be incorporated into the optimization (1) [BOFN18], it is irrelevant to our rest shape optimization since we focus on the static equilibrium case with $\mathbf{\dot{q}}=0$. We can obtain the generalized positions for the next time step by solving the optimization problem (1), e.g., with a gradient-based optimizer (such as gradient descent or Newton’s method [NW06]) using the gradient of the DER objective with respect to $\mathbf{q}$, $\nabla_{\mathbf{q}}E_{\mathrm{DER}}(\mathbf{q})$, while handling the fixed variables as Dirichlet boundary conditions. We note that the gradient with respect to the fixed variables is not defined (or, in practice, it can be assumed 0 for convenience). In our framework, we use a single Newton iteration per simulation step for efficiency [BW98, KE22, HYW23] (although this approach sacrifices stability with larger parameter values [SWP23]). If one prefers avoiding the Hessian of the DER objectives (to make the entire framework Hessian-free), it would be possible to use Alternating Direction Method of Multipliers (ADMM) [NOB16] to solve (1). In the following sections, we define each objective term and its gradient.
@@ 3.1 Inertia
$E_{\mathrm{inertia}}(\mathbf{q})=\frac{1}{2\Delta t^{2}}\left\lVert\mathbf{q}-\mathbf{q}^{*}\right\rVert^{2}_{\mathbf{M}},$ (3)
where $\mathbf{q}^{*}(=\mathbf{q}^{t}+\Delta t\mathbf{\dot{q}}^{t}+\Delta t^{2}\mathbf{M}^{-1}\mathbf{f}_{\mathrm{ext}})$ denotes the predicted generalized positions, $\mathbf{q}^{t}$ and $\mathbf{\dot{q}}^{t}$ denote the generalized positions and velocities at time $t$, respectively, $\Delta t\ (>0)$ time step size, $\mathbf{M}\in\mathbb{R}^{(4N-1)\times(4N-1)}$ a diagonal, generalized mass matrix [BAV10, JNO18], and $\mathbf{f}_{\mathrm{ext}}\in\mathbb{R}^{4N-1}$ a generalized external force vector. While sag-free simulation typically assumes gravity as the sole external force, we can nevertheless easily support other constant external forces, such as wind forces or load on specific vertices [PTC15] or edges.
@@ 3.2 Stretching
where $c_{\mathrm{stretch}}$ denotes the stiffness coefficient for stretching, and $\mathbf{l}_{i}\ (=\left\lVert\mathbf{x}_{i+1}-\mathbf{x}_{i}\right\rVert_{2})$ denotes the length of edge $i$. (Unlike the work of [BAV10], which uses Young’s modulus and shear modulus, we use different coefficients for each energy to individually control their strengths.) Given that the vertices $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ are fixed, there is no stretching energy contribution $E_{\mathrm{stretch},0}(\mathbf{x}_{0},\mathbf{x}_{1})$ from edge 00 (the first edge at the root end).
@@ 3.3 Bending
We define the bending energy on vertex $i$ with the four-dimensional curvature $\bm{\kappa}_{i}\in\mathbb{R}^{4}$ ($\bm{\kappa}_{i}=[\bm{\kappa}_{i,0},\bm{\kappa}_{i,1},\bm{\kappa}_{i,2},\bm{\kappa}_{i,3}]^{T}$, where two 2D curvatures $(\bm{\kappa}_{i,0}$, $\bm{\kappa}_{i,1}$) and ($\bm{\kappa}_{i,2}$, $\bm{\kappa}_{i,3}$) denote the curvature binormals evaluated on the material frames of edges $i-1$ and $i$, respectively, and $\bm{\kappa}=[\bm{\kappa}_{1}^{T},\ldots,\bm{\kappa}_{N-2}^{T}]^{T}\in\mathbb{R}^{4N-8}$) and its corresponding rest curvature $\bm{\bar{\kappa}}_{i}\in\mathbb{R}^{4}$ in 3D [BWR08, FBGZ19] as
@@
where $c_{\mathrm{bend}}$ denotes the stiffness coefficient for bending energy. Here, we do not use the bending model with the two-dimensional curvature (and rest curvature) via the averaged material frames from [BAV10] because their model would fail to correctly evaluate the bending of strands when they are bent or twisted due to the non-unit and non-orthogonal, averaged material frames [GB15, PKLI19]. In addition, the bending model (12) [BWR08, FBGZ19] is slightly simpler than the one presented by [PKLI19] (which modifies the rest length term) and avoids introducing additional complexity when differentiating (12) with respect to the rest length (see Sec. 4.1.3).
@@ 3.4 Twisting
where the twisting energy defined for vertex $i$ is given with twist $\mathbf{m}_{i}$ ($\mathbf{m}=[\mathbf{m}_{1},\ldots,\mathbf{m}_{N-2}]^{T}\in\mathbb{R}^{N-2}$), which takes the reference twist into account [BAV10, KJM10, JNO18, PKLI19], and its corresponding rest twist $\mathbf{\bar{m}}_{i}$ by
$E_{\mathrm{twist},i}(\mathbf{x}_{i-1},\bm{\theta}_{i-1},\mathbf{x}_{i},\bm{\theta}_{i},\mathbf{x}_{i+1})=\frac{1}{2}\left(\frac{c_{\mathrm{twist}}\pi r^{4}}{(\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i})}\right)(\mathbf{m}_{i}-\mathbf{\bar{m}}_{i})^{2},$ (16)
@@
$\nabla E_{\mathrm{twist},i}(\mathbf{x}_{i-1},\bm{\theta}_{i-1},\mathbf{x}_{i},\bm{\theta}_{i},\mathbf{x}_{i+1})=\left(\frac{c_{\mathrm{twist}}\pi r^{4}}{(\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i})}\right)(\mathbf{m}_{i}-\mathbf{\bar{m}}_{i})\nabla\mathbf{m}_{i}.$ (17)
@@ 4 Rest Shape Optimization
We define the rest shape parameter $\mathbf{\bar{s}}=[\mathbf{\bar{l}}_{1},\ldots,\mathbf{\bar{l}}_{N-2},\bm{\bar{\kappa}}^{T},\mathbf{\bar{m}}^{T}]^{T}\in\mathbb{R}^{6N-12}$ since $(6N-12)=(N-2)+4(N-2)+(N-2)$. We aim to optimize $\mathbf{\bar{s}}$ based on minimization of the kinetic energy with a regularizer subject to box constraints. Specifically, we formulate the rest shape optimization as
@@
where $\mathbf{\bar{s}_{\mathrm{min}}}\in\mathbb{R}^{6N-12}$ and $\mathbf{\bar{s}_{\mathrm{max}}}\in\mathbb{R}^{6N-12}$ denote the lower and upper bounds imposed on $\mathbf{\bar{s}}$, respectively, and $F_{\mathrm{kin}}(\mathbf{\bar{s}})$ and $F_{\mathrm{reg}}(\mathbf{\bar{s}})$ denote the kinetic energy objective and a regularizer, respectively. Here, we divide $F_{\mathrm{kin}}(\mathbf{\bar{s}})$ and $F_{\mathrm{reg}}(\mathbf{\bar{s}})$ by $\Delta t^{2}$, anticipating that it will cancel with the $\Delta t^{2}$ in (20) and (29), and noting that this scaling by a constant value has no effect on the rest shape optimization (18). In addition, note that $\mathbf{\bar{s}}$ excludes $\mathbf{\bar{l}}_{0}$ because the first two root-end vertices are fixed, and thus the edge between them should also be fixed. As such, we do not define the Jacobian with respect to $\mathbf{\bar{l}}_{0}$ (similar to the fixed variables in Sec. 3), and thus $\mathbf{\bar{l}}_{0}$ is not modified by the rest shape optimization (18).
While one could consider optimizing $\tilde{F}$ in (18) with respect to $\mathbf{q}$ given its fewer DOFs than those for $\mathbf{\bar{s}}$, and then initializing $\mathbf{\bar{s}}$ (which is eventually used to compute DER objectives and forces) with the optimized $\mathbf{q}$, this approach makes it difficult to control the system energy (e.g., to avoid stability problems) via box constraints. $\mathbf{q}$ is nonlinearly related to multiple components of $\mathbf{\bar{s}}$, and thus it is not straightforward to specify box constraints on $\mathbf{q}$ that enforce, e.g., $|\bm{\bar{\kappa}}_{i,0}|<1$ to regulate bending energy (12). Additionally, computing the Jacobian of the generalized forces with respect to $\mathbf{q}$ (i.e., Hessian) is much more complicated compared to those with respect to $\mathbf{\bar{s}}$ (see Sec. 4.1.2, Sec. 4.1.3, and Sec. 4.1.4).
@@ 4.1 Generalized Kinetic Energy Objective
Considering the Jacobian $\mathbf{J}_{\mathrm{DER}}(=\frac{\partial\mathbf{f}_{\mathrm{DER}}}{\partial\mathbf{\bar{s}}})\in\mathbb{R}^{(4N-8)\times(6N-12)}$ (while treating $\mathbf{M}$ as constant because it is aimed to evaluate the norm in the kinetic energy metric at the initial setting), the gradient of (20) can be computed as
@@
and the Hessian can be approximated in the Gauss-Newton style (ignoring the third-order tensors) by
@@ 4.1.1 Inertia
This result indicates that the inertia force is independent of $\mathbf{\bar{l}}$, and thus the Jacobian $\frac{\partial\mathbf{f}_{\mathrm{inertia}}}{\partial\mathbf{\bar{s}}}=0$. In addition, (23) is also independent of $\Delta t$, and thus the actual value of $\Delta t$ has no effect on the rest shape optimization (18). In the rest shape optimization for static equilibrium, a larger norm of the generalized forces, $\left\lVert\mathbf{f}_{\mathrm{inertia}}\right\rVert_{2}$, typically requires more significant rest shape changes.
@@ 4.2 Regularizer for Rest Shape Parameters
where $\alpha$ denotes a positive regularizer coefficient (we typically set $\alpha=10^{-5}$). Although the given system may already be overdetermined (having a unique solution) depending on the initial strand geometry, even in such cases, using the regularizer generally has a positive effect in mitigating overfitting. The previous work [TKA11] proposed using a regularizer based on the force norm, but we found it to be unnecessary because the regularizer in our framework is used to ensure that the system is overdetermined, while the box constraints can restrict the changes in the rest shape parameters in a more controlled way (see Sec. 4.3).
@@ 4.3 Box Constraints for Rest Shape Parameters
Given the turning angle $\phi_{i}$ at vertex $i$ [BWR08], we can derive the relation between the norm of the curvature and turning angle as $\frac{1}{2}\left\lVert\bm{\kappa}_{i}\right\rVert_{2}^{2}=\left(2\tan\frac{\phi_{i}}{2}\right)^{2}$. Thus, if we have the change of the turning angle $\Delta\phi$ at $\phi_{i}=0$ with the corresponding change of the curvature $\Delta\kappa$ (assuming $\Delta\kappa=\Delta\bm{\kappa}_{i,0}=\Delta\bm{\kappa}_{i,1}=\Delta\bm{\kappa}_{i,2}=\Delta\bm{\kappa}_{i,3}$ for simplicity), we get $\frac{1}{2}(4\Delta\kappa)^{2}=\left(2\tan\frac{\Delta\phi}{2}\right)^{2}$, and thus $|\Delta\kappa|=\sqrt{2}|\tan\frac{\Delta\phi}{2}|$. This relation indicates that, e.g., allowing a maximal change of the turning angle of $\frac{\pi}{2}$ (i.e., $\Delta\phi=\frac{\pi}{2}$) leads to a change of the curvature of $\Delta\kappa=\sqrt{2}$. Based on this analysis for bending, we define $\mathbf{b}$ as
@@ 4.4 Gauss-Newton Solver with Penalty Method
Given the least-squares-style nonlinear objectives arising in our problem, we aim to solve the optimization via the GN method [NW06]. While enforcing box constraints within Newton-type optimizers is possible [TB21], it requires accounting for the box constraints when solving inner problems, which leads to an additional computational cost compared to unconstrained linear solves. Considering that our problems are relatively stiff (albeit yielding sparse systems), instead of handling box constraints in the inner problems, we prefer to convert the box-constrained minimization into an unconstrained one via the penalty method [NW06], thus ensuring unconstrained inner linear systems for greater efficiency.
Specifically, we reformulate the box-constrained minimization (18) into the following unconstrained minimization by converting the box constraints into a penalty objective $F_{\mathrm{box}}(\mathbf{\bar{s}})$ [NW06]:
$\mathbf{\bar{s}}=\operatorname*{arg\,min}_{\mathbf{\bar{s}}}F(\mathbf{\bar{s}}),\ F(\mathbf{\bar{s}})=\frac{1}{\Delta t^{2}}\left(F_{\mathrm{kin}}(\mathbf{\bar{s}})+F_{\mathrm{reg}}(\mathbf{\bar{s}})+F_{\mathrm{box}}(\mathbf{\bar{s}})\right),$ (37)
@@
with $\beta$ denoting a penalty parameter (we set $\beta=10^{6}$). Note that, similar to the original rest shape optimization (18), $\Delta t$ in (37) cancels completely with (20), (29), (38), and therefore $\Delta t$ has no effect on achieving static equilibrium via rest shape optimization.
@@
Algorithm 1 shows our rest shape optimization algorithm with the GN solver. To accelerate its convergence via warm starting and to reduce the possibility of falling into suboptimal local minima, we initialize $\mathbf{\bar{s}}$ from the initial $\mathbf{l}$, $\bm{\kappa}$, and $\mathbf{m}$. The inner linear system is guaranteed to be symmetric positive definite (SPD) due to the GN-style Hessian approximation and the regularizer. Given that the system is relatively stiff yet sparse, we solve it using a sparse Cholesky-based direct solver. To ensure a decrease of the objective, we employ a back-tracking line search [NW06]. In practice, as we enforce the box constraints via the penalty, the rest shape parameters $\mathbf{\bar{s}}$ can slightly violate their bounds. Given these relatively small violations, if necessary, we can project $\mathbf{\bar{s}}$ back into the valid range in each iteration (although this treatment was not needed in any of our examples). We terminate the GN iterations when $\left\lVert\Delta\mathbf{\bar{s}}\right\rVert_{2}$ becomes smaller than a threshold $\epsilon\ (=10^{-5})$. In addition, to ensure that the GN solver ultimately terminates, we halt the iteration if it exceeds $500$ iterations or the back-tracking line search fails.